A 3 kg cart starting from rest on a frictionless surface is pulled by a horizontal massless rope with a constant force of 6 N. What is the acceleration of the cart and how fast is it moving at the end of 2 seconds?

show how to do it and answer please!
1
Expert's answer
2010-11-01T03:46:13-0400
F = ma
a = F/m = 6/3 = 2 m/sec2
v = v0+ at = 0 + 2*2 = 4 m/sec.